OpenStudy (anonymous):

Use implicit differentiation to find an equation of a tangent line to the curve at a given point? x^2 + 2xy - y^2 + x = 2 (1,2) (hyperbola) I got: 2x + 2 (dy/dx) - 2y (dy/dx) = 0 (2 - 2y) dy/dx = -2x dy/dx = -x / (1-y) At (1,2) the slope is: -(1)/(1-(2))=-1/-1=1 y-y1=m(x-x1) (y-2)=1(x-1) (y-2)=1x-1 y-2=1x-1 y = x + 1 can someone verify the correct answer?

OpenStudy (anonymous):

when using implicit differentiation on 2xy, you have to apply the product rule. \[\frac{d}{dx}2xy=2y+2x*\frac{dy}{dx}\]

OpenStudy (anonymous):

Actually the answer is 2y-7x+3=0

OpenStudy (anonymous):

yep he mentioned it you made your mistake at that point.

OpenStudy (anonymous):

can someone show me the steps then.. so I can see what I did wrong?

OpenStudy (anonymous):

you also forgot to apply chain rule on -y^2. \[\frac{d}{dx}-y^2=-2*y*\frac{dy}{dx}\] the whole derivative should look like this \[2x+2y+2xy'-2y*y'+1=0\] I also noticed you didn't differentiate the last x. Try to go back and differntiate everything again and get the rigth answer before moving on to understand it.

OpenStudy (anonymous):

2x+2(xy'(x)+y)-2yy'(x)+1=0 2x+2xy'(x)'+2y-2yy'(x)+1=0 2xy'(x)-2yy'(x)=2x-2y-1 y'(x)(2x-2y)=2x-1 y'(x)=(2x-2y-1)/(2x-2y) (2(1)-2(2)-1)/(2(1)-2(2)) 7/2 y=7/2x-3/2

OpenStudy (anonymous):

That is correct

OpenStudy (anonymous):

yay.. gosh dang lol.